Given geometric Progression 1/64,1/32,1/16,1/8,1/4,… how many terms such that Sn=2^20-1/64 find the last term sn?


1
Expert's answer
2022-05-04T17:34:58-0400
"a=\dfrac{1}{64}, r=2"

"S_n=\dfrac{a(1-r^n)}{1-r}"

"\dfrac{\dfrac{1}{64}(1-2^n)}{1-2}=\dfrac{2^{20}-1}{64}"

"2^n=2^{20}"

"n=20"

"a_{20}=\dfrac{1}{64}(2)^{20-1}=2^{13}=8192"